OpenStudy (anonymous):

Find the solution to the equation 64^(3 – x) = 4^2x Find the solution to the equation 64^(3 – x) = 4^2x @Mathematics

OpenStudy (jamesj):

64 = 4^3 thus \[ 64^{3-x} = 4^{3(3-x)} = 4^{9-3x} \] Now can solve this?

OpenStudy (ash2326):

4^3(3-x)=4^(2x) 9-3x=2x x=1.8

OpenStudy (jamesj):

Now can you solve this? Because now you know the original equation is equivalent to 4^(9-3x) = 4^(2x)

OpenStudy (anonymous):

I got x = 9/5

OpenStudy (jamesj):

right.

OpenStudy (anonymous):

Thank you :)